OpenStudy (anonymous):

Solve by Factoring: 3x2 = -9x What are the solutions for x? x = 3, x = 9 x = 0, x = 3 x = 0, x = -3 none of the above

hero (hero):

3x^2 = -9x 3x^2 + 9x = 0 3x(x-3) = 0 x = 0, 3

OpenStudy (anonymous):

Isn't it -3 and 0?

OpenStudy (anonymous):

3x^2= -9x --- 3x^2+9x=0 3x(x+3) = 0 3x= 0 and x+3=0 x=0 x= -3 x= (0, -3)

hero (hero):

yeah, it's what happens when you write too fast.

OpenStudy (anonymous):

Haha

hero (hero):

x = 0,-3